OpenStudy (anonymous):

Hexagon D and hexagon H are regular hexagons. The scale factor from hexagon D to hexagon H is 0.25. One side of hexagon D measures 18 cm. What is the length of one side of hexagon H?

OpenStudy (mathstudent55):

D/H = 0.25 18/H = 0.25/1 Cross multiply and solve for H

OpenStudy (anonymous):

thanks :)